# Blind Identification of Channel Codes: A Subspace-Coding Approach

### v1

@@ III-B An Improved Decoder for Large NN
Next, computing the subspace distance of $D(Y)$ from a code $\cal C$} involves rank computations of matrices of dimensions at most $(N^{*}+k)\times n$ (by (2)). A rank computation of a matrix of size $a\times b$ requires $\mathcal{O}(ab\min\{a,b\})$ operations. Assuming $N^{*}+k\geq n$, this is $\mathcal{O}((N^{*}+k)^{2}n)$ per subset. Thus, the total rank-computation cost per code is $\mathcal{O}\!\left(l(N^{*}+k)n^{2}\right)$. Combining both, the total complexity of the proposed decoder is $\mathcal{O}\!\left(M\big(Nn(n-k)+l(N^{*}+k)n^{2}\big)\right).$ Expressing $k=Rn$, where $R$ is the code rate, this becomes $\mathcal{O}\!\left(n^{2}\big(MN(1-R)+lN^{*}\big)+MlRn^{3}\right),$ and for fixed $N^{*}$ and $l$, the overall complexity scales as $\mathcal{O}\!\left(MlRn^{3}\right).$
@@ III-C Bound on the Error Probability of Improved $\mathsf{M}\text{-}\mathsf{DenSD}$ Decoder
    Event ${\cal E}_{1}$: For each $R\in{\cal R}$, $\mathsf{rank}(A[{R},:])<k$.
@@
By Theorem 3, the improved $\mathsf{M}\text{-}\mathsf{DenSD}$ decoder will be correct if the complements of both events ${\cal E}_{1}$ and ${\cal E}_{2}$ occur for at least one subset $R$. Thus, the probability of error $\mathsf{P}_{e}$ of the decoder is bounded as follows.
$\mathsf{P}_{e}$ $\leq P({\cal E}_{1})+P({\cal E}_{2})$
Note that $P({\cal E}_{1})=\sum_{r=0}^{k-1}\mathsf{P}(\mathsf{rank}(A[{R},:])=r).$ Now, for a uniformly random binary matrix $A_{1}\in\mathbb{F}_{2}^{N^{*}\times k}$, the rank distribution satisfies $\mathsf{P}(\mathsf{rank}(A_{1})=r)\leq 2^{-(N^{*}-r)(k-r)}$, for $0\leq r\leq\min(N,k)$ (see [36], for arguments based on which we can arrive at this bound).
@@
Observe that the first term in the R.H.S. of both (III-C) and (III-C) does not depend on $N$, and is minimized (attaining a constant floor) as $N^{*}$ increases, for any fixed $N$. On the other hand, the second term is the tail of the binomial distribution with parameters $N,\alpha(p)$, and this increases as $N^{*}$ increases, for any fixed $N$. This indicates a ‘sweet-spot’ for $N^{*}$, for fixed $N$, which was observed via simulations in Subsection III-B. Characterizing this optimizing $N^{*}$ is an ongoing work. As the mean of the binomial distribution is $N\alpha(p)$, we also observe that choosing $N^{*}=cN$ for some constant $c<1-\alpha(p)$, the bound on $\mathsf{P}_{e}$ goes to 0 as $N\to\infty$. $\square$

### v2

@@ III-B An Improved Decoder for Large NN
Next, computing the subspace distance of $D(Y)$ from a code $\cal C$} involves rank computations of matrices of dimensions at most $(N^{*}+k)\times n$ (by (2)). A rank computation of a matrix of size $a\times b$ requires $\mathcal{O}(ab\min\{a,b\})$ operations. Assuming $N^{*}+k\geq n$, this is $\mathcal{O}((N^{*}+k)n^{2})$ per subset. Thus, the total rank-computation cost per code is $\mathcal{O}\!\left(l(N^{*}+k)n^{2}\right)$. Combining both, the total complexity of the proposed decoder is $\mathcal{O}\!\left(M\big(Nn(n-k)+l(N^{*}+k)n^{2}\big)\right).$ Expressing $k=Rn$, where $R$ is the code rate, this becomes $\mathcal{O}\!\left(n^{2}\big(MN(1-R)+lN^{*}\big)+MlRn^{3}\right),$ and for fixed $N^{*}$ and $l$, the overall complexity scales as $\mathcal{O}\!\left(MlRn^{3}\right).$
@@ III-C Bound on the Error Probability of Improved $\mathsf{M}\text{-}\mathsf{DenSD}$ Decoder
    Event ${\cal E}_{1}$: The set ${\cal R}^{\prime}\triangleq\{R\in{\cal R}:\text{wt}_{H}{(E[{i},:])}<\lfloor(\delta-1)/2\rfloor,\forall i\in R\}$ is non-empty, and for every such $R\in{\cal R}^{\prime}$ we have $\mathsf{rank}(A[{R},:])<k$.
@@
By Theorem 3, the improved $\mathsf{M}\text{-}\mathsf{DenSD}$ decoder will be correct if the complement of the event ${\cal E}_{1}\cup{\cal E}_{2}$ occurs Thus, the probability of error $\mathsf{P}_{e}$ of the decoder is bounded as follows.
$\mathsf{P}_{e}$ $\leq P({\cal E}_{1}\cup{\cal E}_{2})\leq P({\cal E}_{1})+P({\cal E}_{2})$
Note that $P({\cal E}_{1})\leq\sum_{r=0}^{k-1}\mathsf{P}(\mathsf{rank}(A[{R},:])=r),$ for any $R\in{\cal R}^{\prime}$ Now, for a uniformly random binary matrix $A_{1}\in\mathbb{F}_{2}^{N^{*}\times k}$, the rank distribution satisfies $\mathsf{P}(\mathsf{rank}(A_{1})=r)\leq 2^{-(N^{*}-r)(k-r)}$, for $0\leq r\leq\min(N,k)$ (see [36], for arguments based on which we can arrive at this bound). Thus, $P({\cal E}_{1})\leq\sum_{r=0}^{k-1}2^{-(N^{*}-r)(k-r)}.$
@@
Observe that the first terms in the R.H.S. of both (III-C) and (III-C) does not depend on $N$, and is decided by the value of $N^{*}$ (given $k$). In particular, this first term in each of (III-C) and (III-C) decreases as $N^{*}$ increases, for any fixed $N$. On the other hand, the second term is the tail of the binomial distribution with parameters $N,\alpha(p)$, and this increases as $N^{*}$ increases, for any fixed $N$. This indicates a ‘sweet-spot’ for $N^{*}$, for fixed $N$, which was observed via simulations in Subsection III-B. Characterizing this optimizing $N^{*}$ is an ongoing work. As the mean of the binomial distribution is $N\alpha(p)$, we also observe that choosing $N^{*}=cN$ for some constant $c<1-\alpha(p)$, the bound on $\mathsf{P}_{e}$ goes to 0 as $N\to\infty$. $\square$